Jarett & Sons's common stock currently trades at $40.00 a share. It is expected to pay an annual dividend of $1.50 a share at the end of the year (D1 = $1.50), and the constant growth rate is 7% a year.

  1. What is the company's cost of common equity if all of its equity comes from retained earnings? Do not round intermediate calculations. Round your answer to two decimal places.

  2. If the company issued new stock, it would incur a 13% flotation cost. What would be the cost of equity from new stock? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

a.Cost of equity=(D1/Current price)+Growth rate

=(1.5/40)+0.07

=10.75%

b.Cost of equity=(D1/Current price(1-Flotation cost)+Growth rate

=(1.5/(40*(1-0.13)+0.07

=(1.5/34.8)+0.07

=11.31%(Approx).